OpenStudy (anonymous):

x-intercept 3, y-intercept 2/3 What is the slope-intercept form from the line?

OpenStudy (anonymous):

X-intercept means (3,0) Y-intercept means (0,2/3) \[Y-0/X-3=(2/3-0)/0-3\]

OpenStudy (anonymous):

\[=2x+9y=6\] \[y=2/9x+2/3\]